Complete the table assuming continuously compounded interest. (Round your answers to two decimal places.)
Initial Investment: UNKNOWN
Annual % Rate: UNKNOWN
Time to Double: 11 Years
Amount after 10 Years: $1800

Answer by htmentor(1343)

The formula for continuously compounded interest is:
A = P*e^(rt), where A = amount, P = principal, r = the interest rate, t = time
Given: Time to double is 11 years; amount after 10 years = 1800
Thus 2P = P*e^(r*11)
Solving for r gives r = ln(2)/11
Now we can find P:
1800 = P*e^(ln(2)/11*10) -> P = 1800/e^(ln(2)/11*10) = $958.54